OpenStudy (anonymous):

Solve the following problem using substitution or elimination: The sum of two numbers is 170 and their difference is 26. Be sure to show all work.

jhonyy9 (jhonyy9):

170/2 =85 85+26/2 =85+13=98 170-98=72 si these are 98 and 72

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

so the two nos are 98 and 72